Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, was - 0:20 Thursday 27 February 2025 (66258) Print this report
ITF recovery after Ncal intervention

We started the recovery around 19:00 UTC.

The interferometer reached CARM_OFFSET_STEP_3_OF_3 fairly quickly, but the CITF loops UGFs were lower than normal and the sidebands power was low as well. We initially thought this to be caused by a drift in TCS power, but Alessio verified and found no issue.

In the meantime Nicola noticed that many of the galvo loops were open, specifically B2_QD2 and B1p_QD1. The former successfully engaged after an open/close cycle, while the latter kept failing because the beam was clearly misaligned on the quadrant.

We contacted Michal to address the situation. With his help, we managed to bring back the SR - SDB1 pair to yesterday's position and we could continue with the lock acquisition.

The ITF reached LN2 twice, but unlocked immediately after engaging BS fbw (Fig.1). It seems that the interplay between the BS fbw loop and the PR AA loop is back.

To patch this issue, we have commented the lines that engage the fbw in ITF_lock.py (5899 e 5900) and engaged the fbw manually at the end of the LN2 timer, but the TX component still caused an unlock.

They are left commented for the night in order to allow the ITF to re-lock on its own. I cannot solve this issue at this time, and I definitely will not be able to tonight.

 

Images attached to this report
Comments to this report:
pinto, mantovani - 10:33 Thursday 27 February 2025 (66263) Print this report

this morning we found the ITF locked with BStx in fbw, BSty in drift control. The BNS range was above 53 Mpc. We didn't want to affect the science data taking phase, so we left the ITF like this and we just uncommentend in ITF_LOCK.py the engagement of BS AA loop.

at the next unlock please reload ITF_LOCK metatron node.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×